------------------------------------------------------------
Date: Sat Aug 5 00:04:09 2000, Shift: 16:00-24:00
Supervisor: JHL, with: TST
Log book#3 pages 142 - 147
Shift Summary:
------------------------------------------------------------
1. Beam
- Cogged beam at gamma = 70 for about an hour.
- Run 2049 has ~310 collision events without trigger6.
(Due to the loose cable connection on the trigger logic: RD fixed it).
- Run 2050 has 280 collision events.
- After the run 2050, No beam due to "Control System Failure".
- Stable beam condition for the weekend is expected, according to MCR.
2. Run 2050
- Time difference in ZDC and BB ~ 0.
(see http://www.rhic.bnl.gov/brahms/WWW/private/list_hyper/brahms-l/0146.html
for the details.)
- trigger6/trigger4 = 23%, trigger6/trigger1=40%
- Si detector seems to work properly: Nice correlation Si ADC vs Tile ADC.
3. Detectors
- All HV Main Frames seem to be happy
- Seen nice tracks in MTPC1 with GG = 180+-80 V.
- T2 is out of DAQ stream due to the readout failure
- Multiple peaks at |TimeLeft-TimeRight| ~ 100 seen for ZDC: Why?
(see http://pii3.brahms.bnl.gov/~operator/pictures/zdc_timediff_2050.gif)
------------------------------------------------------------
This archive was generated by hypermail 2b29 : Sat Aug 05 2000 - 00:07:33 EDT